3. In rabbits, chocolate-colored fur (w is dominant to white fur, straight fur (c is dominant to curly fur (c), and long ear (s is dominant to short ear (s). The cross of a trihybrid rabbit with straight, chocolate-colored fur and long ears to a rabbit that has white, curly fur and short ears produces the results in the table Number Phenotype White, short, straight 13 Chocolate, long, straight 165 Chocolate, long, curly 13 White, long, straight 82 Chocolate, short, straight 436 Chocolate, short, curly 79 162 White, short, curly 450 White, long, curly 1400 a. (1 pt) Determine the order of the genes on the chromosome b. (1 pt) Identify the alleles that are present on each of the homologous chromosomes in the trihybrid rabbits. c. (1 p) Calculate the recombination frequency between c and w pair of genes. d. (1 pt Calculate the recombination frequency between wand s pair of genes. e. (1 pt Determine the interference value for this cross.
Answer #1

Answer:

a).

The order of genes is -----

s------25.21 m.u.---------- w---13. 36m.u.--c

b). s w+ c+ / s+ w c

c). RF between c & w = 13.36%

d). RF between w & s = 25.21%

e). 0.45

Explanation:

Shorthand

Number of progeny

Class

w s c+

13

double cross over

w+ s+ c+

165

double cross over

w+ s+ c

13

single cross over

w s+ c+

82

single cross over

w+ s c+

436

single cross over

w+ s c

79

single cross over

w s c

162

Parental cross over

w s+ c

450

Parental cross over

Total

1400

Imagine the parental genotype of gametes are w+ s c+ & w s+ c

1. If single cross over (SCO) occurs between w+ & s and w & s+

Normal order= w+---------s & w-----s+

After cross over= w+------s+ & w-----s

w+------s+ recombinants are 165+ 13 = 178

w-----s recombinants are 162 + 13 = 175

Total recombinants = 353

RF = (353/1400.)*100 =25.21%

2. If single cross over (SCO) occurs between s & c+ and s+ c

Normal order= s---------c+ & s+----c

After cross over= s-----c & s+------c+

s-----c recombinants are 79+162= 241

s+------c+ recombinants are 165+82 = 247

Total recombinants = 488

RF = (488/1400)*100 = 34.86%

3. If single cross over (SCO) occurs between w+ & c+ and w & c

Normal order= w+---------c+ & w------c

After cross over= w+-----c & w------c+

w+ -----c recombinants are 13+79= 92

w------c+ recombinants are 13+82= 95

Total recombinants = 187

RF = (187/1400)*100 =13.36%

% RF = Map unit distance

Distance between the w & s = 25.21 mu

Distance between the s & c = 34.86 mu

Distance between the w & c = 13.36 mu

The order of genes is -----

s------25.21 m.u.---------- w---13. 36m.u.--c

RF between s & w = 25.21% = 0.2521

RF between w & c = 13.36% = 0.1336

Expected double cross overs = 0.2521 * 0.1336 = 0.0337

The correct parental genotype of gametes are s w+ c+ & s+ w c

Observed double cross overs are s w c+ & s+ w+ c = 13 + 13

= 26/1400 = 0.0186

Observed double cross progeny are more than the expected double cross over progeny.

Coefficent of coincidence (CC) = Obeserved dco / Expected dco

CC = 0.014/0.032 = 0.55

Interference = 1-CC = 1-0.55 = 0.45